The perihelion and aphelion distances for Mercury are 45.9 x 10^9 m and 69.8 x 10^9 m respectively. The speed of Mercury at aphelion is 3.88 x 10^4 m/s. What is the speed at perihelion.
Answer #1

Ven o Cmservabion ot anlar maweubun for Paneh A F dp 45.9 Xlo
Answer #2

SOLUTION :


Let :


Ra = distance of aphelion for Mercury = 69.8 * 10^9 m


Va = speed of Mercury at aphelion = 3.88 * 10^4  m/s 


Rp = distance of perihelion for Mercury = 45.9 * 10^9 m


Vp = speed of Mercury at aphelion = To be found 


Angular speed remains same at both points.


So,


Va / Ra = Vp / Rp


So,


Vp 

= Va Rp / Ra 

= 3.88 * 10^4 * 45.9 * 10^9 / (69.8 * 10^9)

= 2.55 * 10^4 m/s (ANSWER)

answered by: Tulsiram Garg